Mr. David B. Carpenter 98 Gordon Avenue Portswood Southampton S02 1DA

July 28, 1992

Woodworm Records Ltd, P.O. Box 37 Banbury Oxon OX16 8YN

Dear Woodworm Records,

I seem to have ordered the wrong ticket for this year’s Cropredy festival. I intended to order a full weekend ticket, but I must have specified a Saturday-only ticket by mistake.

I hope you will be able to exchange this ticket. I’m including the ticket, together with a cheque to cover the difference in price (£27 − £18 = £9). Apologies for the inconvenience,

Yours sincerely,